1. The total cost, C(x) associated with producing and marketing x units of an item is given by

C (x) = 0.005x3 -0.02x2 -30x +3000

Find (i) Total cost when output is 4 units.

(ii) Average cost of output of 10 units.

(iii) Marginal cost when output is 3 units.




1
Expert's answer
2021-03-23T11:21:04-0400

C=0.005X3-0.02X2-30X+3000

i) Total Cost when the output is 4 units

"TC=0.005(4)"3"-0.02(4)"2"-30X+3000"

"TC=0.005(64)-0.02(16)-120+3000"

"TC=0.32-0.32-120+3000"

"TC=2,880"

ii) Average cost is equal to total cost (TC) divided by the number of units of a good produced.

Average Cost (AC) = TC/Q

AC = (0.005X3-0.02X2-30X+3000)/X

"AC=0.005X"2"-0.02X-30+3000\/X"

"AC=0.005(10)"2"-0.02(10)""-30+3000\/10"

"AC=0.5-0.2-30+300"

"AC" "=270.30"

iii) Marginal Cost represents the incremental costs incurred when producing additional units of a good or service. It is calculated by taking the total change in the cost of producing more goods and dividing that by the change in the number of goods produced.

"MC=dTC\/dX=0.015X"2"-0.04X-30"

"MC=0.015(3)"2"-0.04(3)-30"

"MC=0.135-0.12-30"

"MC=""-29.985"